OpenStudy (anonymous):

One number is 8 times another number. The numbers are both positive and have a difference of 70. Find the numbers.

OpenStudy (agreene):

8x=y y-70=x --------------------- 8x-70=x 7x=70 x=10 ============ 10*8=y 80=y

OpenStudy (anonymous):

thanks